Please provide an easy-to-understand explanation.
Thank you.

My father is 47 years old today and my mother is 38. When they were married, my father was exactly half as old again as my mother. How long ago was their wedding?

Answer by ankor@dixie-net.com(22740)

My father is 47 years old today and my mother is 38.
When they were married, my father was exactly half as old again as my mother.
How long ago was their wedding?
:
let t = no. of years from their wedding
47 - t = 1.5(38-t)
47 - t = 57 - 1.5t
-t + 1.5t = 57 - 47
.5t = 10
t = 20 yrs since the wedding
:
:
Check this:
father was 27, mother was 18, when married
1.5 * 18 = 27
: